OpenStudy (anonymous):

Describe the behavior of f(n)=n^(1/n) in words.

OpenStudy (anonymous):

End behavior? or as n approaches 0? or both?

OpenStudy (anonymous):

In any case, you'll have to take the limit: \[\large \lim_{n\to c}f(n)=\cdots\] where you choose \(c=0,\pm\infty\).

OpenStudy (anonymous):

My paper says as n grows large

OpenStudy (anonymous):

End behavior it is! Need help finding the limits? : \[\lim_{n\to\infty}n^{1/n}\] \[\lim_{n\to-\infty}n^{1/n}\]

OpenStudy (anonymous):

The reason I can't figure it out is because at one point in an example question, it says f(n) approaches a limit of 2.7 because the graph continually increases, but never gets equal to or greater than 2.7. The graph f(n)=\[n ^{1/n}\] increases to a certain point and then decreases, so what am I supposed to say?

OpenStudy (anonymous):

Well it sounds like, in addition to finding the limit at infinity, you also have to find the local maximum for positive \(n\). Is that right? To find the limit: do you know L'Hopital's rule?

OpenStudy (anonymous):

I don't think so:( I took precalculus last year and this is the assignment to start us into calculus. What is L'Hopital's rule?

OpenStudy (anonymous):

You'll learn soon enough. It's a fairly easy way to determine limits at 0 and ±infinity. To find this limit, I think it would suffice to show what happens for certain values of \(n\). For example, see the following table: The value of \(n^{1/n}\) increases for a while then dips back down for some \(n\) between 2 and 4. To find out exactly when the dip occurs, you need to use the first derivative test. Have you learned that yet?